3. Given the following advertising expenditure - sale volume data, Find sample mean of X sample mean of Y, sample variance of X, sample variance of Y, sample Covariance of X and Y, and sample correlation between X, and Y adv exp sales vol sales region 89 87 98 110 103 114 116 110 126 130 5 8 4 5 10 12 13 14 9 10
Answer #1

86 74 25 2 2 2 1 4 6 9 5 620026 00000000003 97803460 88910111 00000000005 00000000009 5678901234 01234567890

from above sample mean of X =9.5

sample mean of Y =108.3

sample variance of X =SSx/(n-1) =82.5/(10-1)=9.1667

sample variance of Y =SSy/(n-1) =1842.1/(10-1)=204.68

sample covariance =Sxy/(n-1)= 365.5/(10-1) =40.6111

correlation coefficient r= Sxy/(√Sxx*Syy)               = 0.9376